Find three consecutive numbers whose sum is 60.

Answer by CharStar(110)

1st Number = x
2nd Number = x+1
3rd Number = x+2
x + x + 1 + x + 2 = 60
3x + 3 = 60
Subract 3 from both sides
3x = 57
x=19
x(19) + 1 = 20
x(19) + 2 = 21
19 + 20 + 21 = 60
Hope that helps you